OpenStudy (anonymous):

1+cos(12x) = ____ a. 2sin^2(12x) b. 2cos^2(6x) c. 2sin^2(6x) d. 2cos^2(12x)

OpenStudy (raden):

b

OpenStudy (anonymous):

why is it b??

OpenStudy (raden):

look the folowing identity : cos^2 x = (1+cos2x)/2 or 1+cos2x = 2*cos^2 x now, if u convert "x" be "6x", it will be 1+cos2(6x) = 2*cos^2 6x, right ? it similar with 1+cos12x = 2*cos^2 6x

OpenStudy (raden):

r u there @Spectrum

OpenStudy (anonymous):

yeah ..ty